CAM, a.a. 2003-2004 - Esercizi 5

Giampiero Palatucci 24-26 marzo 2004

1. Calcolare i seguenti integrali di funzioni razionali:

a. ∫ \frac{x^{2}}{(1 − x)(x − 2)}dx; b. ∫ \frac{2x − 5}{(x^{2} + 2x + 1)}dx; c. ∫ \frac{dx}{x^{3} + x};

d. ∫ \frac{5x + 1}{(x − 1)(x^{2} + 2)}dx; e. ∫ \frac{x^{2} − 1}{(x − 2)(x^{2} + 1)}dx; f. ∫ \frac{x^{2} − 1}{x^{2}(x^{2} + 1)}dx;

g. ∫ \frac{3x − 2}{(x − 1)(x^{2} + 2x + 2)}dx; h. ∫ \frac{x − 5}{(x^{2} + x + 1)}dx;

2. Calcolare i seguenti integrali (per sostituzione): a. ∫ \frac{1 + cos x}{1 − cos x}dx; b. ∫ \frac{1 + cos x}{sin x}dx; c. ∫ \frac{sin x}{4 − 5 sin x}dx;

d. ∫ \frac{1}{sin 2x + cos^{2} x}dx; e. ∫ \frac{1 − sin^{2} x}{1 − 2 sin^{2} x}dx; f. ∫ \frac{2 − sin^{2} x}{sin 2x}dx;